### v1

# SciClaims: An End-to-End Generative System for Biomedical Claim Analysis

Abstract
Validating key claims in scientific literature, particularly in biomedical research, is essential for ensuring accuracy and advancing knowledge. This process is critical in sectors like the pharmaceutical industry, where rapid scientific progress requires automation and deep domain expertise. However, current solutions have significant limitations. They lack end-to-end pipelines encompassing all claim extraction, evidence retrieval, and verification steps; rely on complex NLP and information retrieval pipelines prone to multiple failure points; and often fail to provide clear, user-friendly justifications for claim verification outcomes. To address these challenges, we introduce SciClaims, an advanced system powered by state-of-the-art large language models (LLMs) that seamlessly integrates the entire scientific claim analysis process. SciClaims outperforms previous approaches in both claim extraction and verification without requiring additional fine-tuning, setting a new benchmark for automated scientific claim analysis.
SciClaims: An End-to-End Generative System for Biomedical Claim Analysis
Raúl Ortega Language Technology Research Lab Expert.ai / Madrid, Spain rortega@expert.ai                        José Manuel Gómez-Pérez Language Technology Research Lab Expert.ai / Madrid, Spain jmgomez@expert.ai
1 Introduction
Systematic Literature Review (SLR) is a crucial field of activity in the biomedical and pharmaceutical industries, as well as in hospital settings. It helps identify clinical data on treatment safety and performance, assess the state of the art, and support regulatory and clinical decision-making. In the pharmaceutical industry, SLR supports activities such as R&D, business development, and medical affairs, while in hospitals, it confirms best practices and guides clinical decisions. By enabling access to relevant information, SLR also shortens regulatory review cycles and facilitates the faster market entry of better treatments.
A key task in SLR is to verify that claims made in scientific texts are supported by previously published, peer-reviewed statements. This process helps define a clear information perimeter for a given work, increasing confidence in the scientific literature. Beyond its role in validating research, the verification of scientific claims also plays a crucial role in key business areas such as detecting patent infringement. In the pharmaceutical industry, for example, companies need to analyze prior scientific publications and patents to defend intellectual property rights. This is particularly relevant in high-stakes legal disputes, such as the recent litigation between Pfizer and BioNTech against Moderna over the mRNA COVID-19 vaccine. [^1] https://iclg.com/news/22358-moderna-wins-german-covid-19-patent-infringement-battle
In this work, we address some of the main challenges that prevent scientific claim analysis systems from becoming a practical solution in such competitive markets as SLR and the pharmaceutical industry, including low accuracies, limited throughput, multi-staged complex pipelines with multiple possible failure points, and lack of user-friendly explanations. In doing so, we look into the different steps involved in the scientific claim analysis pipeline, ranging from the extraction and generation of claims from the original text to the retrieval of potentially relevant evidence to validate those claims and the actual verification against the evidence, supporting or refuting them. Based on state-of-the-art large language models (LLMs), our system SciClaims [^2] SciClaims demo (user/pass: guest/aclindustry2025) available at https://labdemos.expertcustomers.ai/health_claims [^3] Code at: https://github.com/anonymous-sciclaims approaches the entire scientific claim analysis pipeline end-to-end, outperforming previous state-of-the-art methods in both scientific claim extraction and verification without requiring additional fine-tuning.
Figure 1: System Architecture.
2 Related Work
The task of analyzing scientific claims from real-world texts based on background knowledge consists of three primary components: claim extraction, evidence retrieval from a document corpora, and verifying or fact-checking the claims against the evidence Eldifrawi et al. (2024); Vladika and Matthes (2023). However, solving this entire pipeline in an end-to-end manner across the three stages is a challenge that still needs to be met.
Several studies have focused on the challenge of claim extraction from texts. Some frameworks based on zero-shot models Pan et al. (2021); Wright et al. (2022) have achieved promising results, focusing on the generation of datasets of claims from raw texts to train fact-checking models on specific domains. Recent approaches based on LLMs employ these models to extract atomic facts Min et al. (2023); Chern et al. (2023), which serve as summaries of the texts from which they are derived.
For the evidence retrieval phase, dense passage retrieval methods, such as ColBERT Khattab and Zaharia (2020), have emerged due to their ability to retrieve highly relevant documents from large corpora with great precision. However, these methods are computationally expensive and therefore less practical for real-time, lightweight applications. Thus, simpler approaches such as BM-25Robertson and Zaragoza (2009) and Elasticsearch, known for their balance between retrieval quality and computational efficiency, are still preferred for deployment in production environments.
The claim verification component of the pipeline has seen notable advancements as well. The emergence of several claim analysis datasets has resulted in numerous frameworks and models that try to solve this challenge. VERT5ERINI Pradeep et al. (2021), PARAGRAPHJOINT Li et al. (2021), and MultiVerS Wadden et al. (2022) have shown promising results in scientific claim verification datasets such as SciFact Wadden et al. (2020) or CLIMATE-FEVER Diggelmann et al. (2021). Furthermore, due to the recent progress in text generation using large language models (LLMs), a significant body of work has been dedicated to the verification of these artificially generated texts. Frameworks such as FactScore Min et al. (2023), FacTool Chern et al. (2023), or LLM Oasis Scirè et al. (2024) focus on fact-checking automatically generated texts rather than existing human-written passages. However, by leveraging their underlying mechanisms, such as the extraction of atomic knowledge units and claim verification through LLMs, these approaches can also be extended to process scientific documents.
Although individual components have seen substantial progress, an integrated system capable of seamlessly connecting these steps remains an open problem. Many existing systems, such as FactDetect Jafari and Allan (2024) or CliVER Liu et al. (2024) focus on claims that are already identified, neglecting the crucial first step of extracting relevant claims from raw, real-world texts. As a result, these systems are limited to pre-identified claims rather than addressing the full pipeline of claim extraction, retrieval, and verification.
Given the complexity of this problem, our approach seeks to optimize each stage of the pipeline for practical deployment. In particular, we aim to refine the claim extraction, retrieval, and verification modules in a manner that ensures both efficiency and accuracy for real-world applications. By drawing on the strengths of existing systems and focusing on improving their integration, we propose a comprehensive solution for the analysis of biomedical claims in real world texts.
3 Architecture
We present a unified and lightweight architecture built around a Llama3 7B model et al. (2024) and an Elasticsearch index with a verification dataset containing peer-reviewed health and biomedicine literature. This architecture is designed to be lightweight enough to run efficiently on a single 24GB VRAM GPU, enabling high-throughput processing through inference frameworks such as vLLM Kwon et al. (2023), which provide exceptional computation speeds. A seamless processing workflow (see Figure 1) is supported, from claim extraction to evidence retrieval and verification. Rationales in natural language are also provided, which justify system outputs. Next, we present the main building blocks of SciClaims.
Claim extraction: This first module extracts potential claims from the source text. A claim is characterized by a specific set of properties detailed in section 5. The Claim Extraction module calls the Llama3 7B model twice: first, to generate an initial list of claims, and second, to refine and filter them, improving the quality of the resulting claims. All the prompts used in our pipeline are provided in Appendix A. Based on our evaluation, we made adjustments to the initial prompts to improve claim extraction performance. Further details on this refinement can be found in section 5.
Document retrieval: The second module retrieves potentially relevant documents from the verification dataset, using the claim as a query to the Elasticsearch index. The verification dataset contains 4.7 million abstracts from PubMed (2000–2022) that were curated using theSemantic Scholar’s Highly Influential Citations metric444www.semanticscholar.org/faq/influential-citations, ensuring that each article is backed by at least three highly influential citations. This selection criterion helps prioritize documents that have been extensively referenced in the academic community, enhancing the quality and relevance of retrieved information for verification. We chose not to filter documents based on Elasticsearch’s scoring mechanism to maximize recall. Instead, the subsequent verification module is responsible for discarding irrelevant documents.
Claim verification: The final module performs fact-checking by making another call to the LLM, providing the claim along with the retrieved related documents. The model assigns one of the following three labels: 1) SUPPORT if the claim is verified by the document, 2) REFUTE if the claim is refuted by the document, or 3) NEI (not enough information) if the document lacks sufficient evidence or is not relevant to the claim. To improve transparency and interpretability, we also request the model to provide a rationale for its decision, including identifying the most relevant sentence(s) that support its conclusion.
4 System implementation
In this section, we demonstrate the functionality of our approach through a web-based claim analysis tool that allows users to easily interact with the system and analyze claims within relevant texts, leveraging the backend architecture presented in the previous section. The user interface provides a drop-down menu featuring over 30 pre-selected examples from various domains, including biomedicine papers, COVID-related news, social media, and patents (see Appendix B). These examples were chosen to represent a broad spectrum of platforms and relevant disciplines, allowing users to explore diverse contexts. By selecting any example from the list, the corresponding text will be displayed in the text box below, with a hyperlink to the original source. The text box is fully editable, enabling users to modify the content and analyze their own text. While the application can process texts up to 10,000 characters, we recommend keeping the input under 2,000 characters for faster results.
As shown in Appendix B, once the user has selected or entered a text, they can click the Analyze button to initiate the claim analysis process. Results are presented as a list of identified claims, each of which can be expanded to reveal detailed information. For each claim, the interface shows:
  • Prediction label: Whether the claim is supported (green) or refuted (red). Results labeled as Not Enough Information by the LLM are not be presented to the user.
  • Prediction score: A normalized probability score that represents the confidence level of the model in its prediction. This score is derived from the statistical outputs of the model, particularly the tokens representing the label string, and it is expressed as a percentage.
  • Evidence: The related document selected by the retrieval module, along with its DOI. The specific sentence(s) in the abstract of the paper that were most influential for the prediction according to the LLM are highlighted.
  • Rationale: A justification of the reasoning behind the classification, providing additional insight into the decision-making process carried out by the model.
5 Experimentation and Results
We evaluate our system following a two-staged approach. First, we compare the individual modules on different standard datasets. Next, we take a more holistic approach and evaluate the performance of the system at every step of the pipeline using a large language model (LLM) as a judge.
5.1 Evaluation with datasets
We evaluate the two main modules of the system (claim extraction and claim verification) against curated datasets to assess the performance of the LLMs for each task in isolation.
5.1.1 Claim extraction
For the evaluation of the claim extraction module, we selected three datasets: SciFact Wadden et al. (2020), FEVER Thorne et al. (2018), and RoSE Liu et al. (2023). SciFact consists of claims derived from scientific articles, paired with evidence from the source document. We use manually crafted claims and their source paragraphs as ground truth to evaluate factual claim extraction. FEVER, which contains general knowledge claims with Wikipedia-based evidence, helps assess performance on non-specialized texts, using Wikipedia passages as ground truth. Since these datasets highlight only selected portions of source content, we compare baseline-generated claims to ground truth using the Levenshtein distance metric, selecting pairs with a similarity score of 0.3 or higher, where such threshold was empirically determined. Additionally, RoSE, originally designed for summarization, helps us to evaluate how well our system synthesizes source abstracts into structured, factually accurate claims.
Method FEVER RoSE SciFact Sentence tokenizer 0.6549 0.6667 0.7211 Pan et al 0.7147 0.7128 0.6561 Wright et al 0.7023 0.6786 0.6475 Noun Phrases Gen 0.7018 0.6889 0.6693 SciClaims 0.7028 0.7184 0.725
Table 1: Similarity Scores in Claim Generation
As a first baseline, we select a sentence tokenizer, treating each sentence in the source paragraph as a claim. Next, we select two baselines Pan et al. (2021); Wright et al. (2022) that use a pipeline with two transformer models to generate claims from the source paragraph through entity extraction. The first model generates a question-answer pair, where the answer is the entity, while the second reformulates it as a claim. We also introduce a method where we propose to build the claims around noun phrases rather than named entities.
To assess the performance of the claim generation methods we compute a similarity score using a DeBERTa model fine-tuned on the STS-B dataset [^5] https://huggingface.co/bobox/DeBERTa2-0.9B-ST-stsb. We compare the most similar generated claim with each ground truth claim extracted from the same source paragraph. If there is not any ground truth claim that surpasses our Levenshtein threshold for a given generated claim, the claim is not taken into account. Table 1 shows that SciClaims achieves the highest score on SciFact and RoSE, outperforming the baselines. However, SciClaims is slightly outperformed by Pan et al. (2021) in FEVER, suggesting that our prompt may need additional refinement to generalize to broader domains. Appendix C reports more extensive results.
5.1.2 Claim verification
The verification task involves predicting the veracity of claim-evidence pairs, classifying each as SUPPORT, REFUTE, or NOT ENOUGH INFORMATION (NEI). Thus, this module is evaluated using the FEVER Thorne et al. (2018) and SciFact Wadden et al. (2020) datasets, where annotated claims-evidence pairs serve as ground truth. Additionally, we use CLIMATE-FEVER Diggelmann et al. (2021), a domain-specific dataset focusing on climate science.
Architecture FEVER CLIMATE-FEVER SciFact Roberta-base* 0.9147 0.4549 0.522 MultiVerS* 0.9381 0.6371 0.7303 SciClaims 0.6467 0.5963 0.6788
Table 2: F1-Scores in Claim Verification. RoBERTa-base and MultiVerS are fine-tuned models, while SciClaims is zero-shot.
To evaluate claim verification we compare our system against two baselines: A fine-tuned RoBERTa-base model Liu et al. (2019) and MultiVerS Wadden et al. (2022), each trained on the respective objective dataset. We evaluate Scifact on its test set, FEVER on its "labelled dev" subset, and we randomly split the CLIMATE-FEVER dataset to follow a standard subset distribution (80% train, 10% validation and 10% test). On these datasets, the fine-tuned models outperform our zero-shot, LLM-based system SciClaims, especially on the FEVER dataset (see Table 2). However, the performance gap significantly narrows on domain-specific datasets like SciFact and CLIMATE-FEVER, highlighting the potential of SciClaims in specialized domains without requiring task-specific training.
5.2 Evaluation with a judge model
The second stage of our evaluation provides a more comprehensive assessment of the entire system using a large language model as judge. Based on the Judge Arena leaderboard [^6] https://huggingface.co/spaces/AtlaAI/judge-arena, we select Qwen 2.5 72B turbo as judge, since it is the first ranked model with open weights and a higher parameter count than our system’s LLM (Llama3 7B). We chose its 4-bit quantization version (Qwen2.5 72B AWQ777huggingface.co/Qwen/Qwen2.5-72B-Instruct-AWQ) due to hardware limitations. For the evaluation sample, we randomly select 120 documents from the PubMed dataset presented in section 3. The evaluation is conducted in three phases: Claim quality, document retrieval, and claim verification and full-system evaluation.
5.2.1 Claim quality evaluation
In this phase, we evaluate the quality of the claims generated by each method. We devise a questionnaire with eight yes/no questions (see Table 3) to capture the desired properties in a correct claim and ask the judge model to answer it. The first question requires context from the source paragraph, while the remaining questions focus solely on the claim. Correct claims need to receive a Yes to all questions. Table 4 shows that our LLM-based system outperforms all other methods, scoring 15 points higher than the second-best, the sentence tokenizer.
Id Question
Q1 Is the claim grounded by the original text?
Q2 Is the claim grammatically correct?
Q3 Does the claim have all the necessary components (subject, predicate, and relevant qualifiers) to form a complete thought?
Q4 Is the claim precise and specific rather than vague?
Q5 Does the claim introduce new information rather than just restating common knowledge?
Q6 Is the claim concise without losing essential information?
Q7 Does the claim provide enough information to be understood independently?
Q8 Would verifying the claim add value to public knowledge?
Table 3: Questions asked to the judge to evaluate the quality of the generated claims.
Method Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Correct Claims (%) Candidate claims
Sentence tokenizer 0.9987 0.9100 0.8436 0.8709 0.9322 0.9009 0.3703 0.9596 32.46 767
Pan et al 0.2449 0.1303 0.1618 0.1371 0.2112 0.1169 0.0337 0.1641 3.15 445
Wright et al 0.8043 0.4714 0.5040 0.3842 0.5593 0.4829 0.1641 0.6249 11.28 4175
Noun Phrases Gen 0.3001 0.1562 0.1857 0.1209 0.2388 0.1320 0.0316 0.2203 2.42 4335
SciClaims 0.9949 0.9474 0.9718 0.9089 0.9320 0.9345 0.5238 0.9756 47.37 779
SciClaims+CDP 0.9942 0.9690 0.9845 0.9380 0.9574 0.9593 0.5930 0.9903 55.43 516
SciClaims+CDP+CR 0.9923 0.9787 0.9884 0.9845 0.9806 0.9748 0.7810 0.9903 76.36 516
Table 4: Claim extraction (phase 1) evaluation results with a judge model (Qwen2.5 72B AWQ).
System All Claims Correct Claims
Generation Module Verification Module Label Accuracy Not NEI (%) Label Accuracy Not NEI (%) System Score Time/doc (secs.)
Sentence tokenizer MultiVerS 0.5494 4.77 0.5127 4.82 0.1664 2.36
Pan et al MultiVerS 0.7591 2.31 0.5714 11.90 0.0180 2.17
Wright et al MultiVerS 0.5015 5.53 0.4268 8.35 0.0482 14.25
Noun Phrases Gen MultiVerS 0.5965 3.08 0.4825 6.67 0.0117 15.01
SciClaims MultiVerS 0.5944 6.02 0.5709 6.50 0.2704 2.86
SciClaims+CDP MultiVerS 0.5867 5.71 0.5361 5.60 0.2971 1.82
SciClaims+CDP+CR MultiVerS 0.6204 3.72 0.5922 7.44 0.4522 2.93
Sentence tokenizer SciClaims 0.6041 61.20 0.6693 59.17 0.2173 11.72
Pan et al SciClaims 0.6568 39.60 0.619 71.43 0.0195 7.61
Wright et al SciClaims 0.6397 69.44 0.7098 74.66 0.0801 65.28
Noun Phrases Gen SciClaims 0.6608 53.02 0.6762 70.48 0.0164 68.01
SciClaims SciClaims 0.6361 58.49 0.6567 59.71 0.3111 12.38
SciClaims+CDP SciClaims 0.6296 58.87 0.6364 62.47 0.3527 8.13
SciClaims+CDP+CR SciClaims 0.6589 53.06 0.6574 54.23 0.5020 9.24
Table 5: Verification and full-system evaluation (phase 3) results with a judge model (Qwen2.5 72B AWQ). The processing time per document is the average time in seconds that each system needs to process a single document.
To further enhance the results, we refined our claim extraction prompts with two optimizations:
  • Claim Definition Properties (CDP). Here we enhance the prompt by incorporating the characteristics of a claim, such as precision, conciseness, or check-worthiness. These characteristics are derived from the questionnaire used to evaluate the quality of the claims. The goal is to guide the LLM to adhere to these criteria when generating the list of claims.
  • Claim Refinement (CR): This upgrade involves a follow-up call to the LLM in order to refine the initial list of claims. For each candidate claim, we pair it with the source paragraph and ask the LLM to refine the claim based on the same criteria (precision, conciseness, check-worthiness, etc.). This step aims to eliminate poorly formed claims and reinforce the application of the specified criteria.
These two upgrades result in an additional 29-point increase in the percentage of correct claims generated by SciClaims while reducing the number of candidate claims generated by our approach. Notably, QA-oriented baselines, such as Wright et al. (2022) and our noun phrase-based generation method, generate significantly larger quantities of claims compared to the other methods in Table 4.
5.2.2 Document retrieval evaluation
We assess document relevance by asking the judge model whether each retrieved paragraph aids claim verification. Recall is evaluated at k = 1,3,5 retrieved documents per claim. As shown in Table 6, claims generated by Wright et al. (2022) retrieve more potentially relevant documents than SciClaims. However, SciClaims performs well, retrieving at least one relevant document in 75% of cases when fetching five documents per claim, a common real-world scenario. This highlights its balance between claim accuracy and document relevance. QA-oriented methods generate more compact, less specific claims than LLMs, increasing the likelihood of retrieving related documents.
Claims All Claims Correct Claims R@1 R@3 R@5 R@1 R@3 R@5 Sentence tokenizer 0.5591 0.7205 0.7795 0.6265 0.7390 0.7871 Pan et al 0.2970 0.4653 0.5248 0.6429 0.7143 0.7143 Wright et al 0.5860 0.7474 0.8038 0.7113 0.8365 0.8726 Noun Phrases Gen 0.4634 0.6456 0.7013 0.6286 0.7619 0.8095 SciClaims 0.5045 0.6645 0.7135 0.5257 0.6938 0.7344 SciClaims+CDP 0.5146 0.6940 0.7466 0.5944 0.7448 0.7867 SciClaims+CDP+CR 0.4727 0.6250 0.6777 0.5102 0.6675 0.7107
Table 6: Document retrieval (phase 2) evaluation results with a judge model (Qwen2.5 72B AWQ).
5.2.3 Claim verification and full-system
SciClaims demonstrates superior claim verification accuracy, outperforming the fine-tuned MultiVerS Wadden et al. (2022) model despite operating in a zero-shot setting. As shown in Table 5, SciClaims consistently produces more accurate labels across all claim generation methods, except Pan et al. (2021). Notably, SciClaims returns significantly fewer NEI labels than MultiVerS, providing greater value to users by delivering more definitive SUPPORT or REFUTE labels. The unusually high accuracy of Pan et al. likely stems from its excessive NEI labels, which simplify label selection. Furthermore, the SciClaims+CDP+CR system achieves the highest overall performance, correctly labeling 50% of generated claims, outperforming the next-best MultiVerS-based system by over five points.
6 Conclusion
Built on state-of-the-art LLMs, SciClaims streamlines claim extraction, evidence retrieval, and verification, outperforming previous methods in both claim extraction and verification without additional fine-tuning. It also delivers user-friendly explanations in natural language, ensures competitive throughput in terms of processing time per input document, and leverages its underlying LLM’s context window to handle larger documents than previous encoder-based approaches. Moreover, SciClaims is also frugal hardware-wise. Major pharmaceutical companies like Sanofi and scientific literature review platforms such as DistillerSR are currently exploring SciClaims for real-world applications including patent infringement detection.
Appendix A Prompts
A.1 SciClaims claim extraction first step
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
Your task is to generate a list with the main factual claims stated in a text. A factual claim makes an assertion about something regarding the subject matter that can be proved or contradicted with factual evidence. Factual claims must be expressed as meaningful, self-contained sentences. Do not include narrative context and disregard absolutely ALL self-referential parts.
Arrange your output using the format:
claim
claim
claim.
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
TEXT: {text}
A.2 SciClaims+CDP
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
Identify and list the main scientific claims stated in a passage. Each claim must satisfy the following criteria:
- **Convey an insight, interpretation, or conclusion drawn from the passage that is testable and generalizable**: The claim should assert an outcome, capability, or effect rather than merely describing a method, aim, or process.
- Example (Good): "Neural networks outperform decision trees in image classification tasks." (Testable outcome)
- Example (Bad): "The proposed method aims to use neural networks for better image classification." (Descriptive, not assertive)
- **Be expressed as a meaningful, self-contained statement**: Each claim should be fully understandable on its own, without needing context from the passage or other claims. It must convey a complete, independent idea. If referencing a study, survey, result, or process, phrase it as a general, verifiable claim.
- Example (Good): "The Amazon rainforest is home to over 10 million species."
- Example (Bad): "As mentioned earlier, the Amazon is one of the most biodiverse places in the world." (Requires prior context and doesnt stand alone.)
- **Emphasize generalization and scientific assertion**: Avoid descriptive or narrative conclusions.
- Example (Good): "Exposure to blue light before sleep can reduce melatonin production." (Generalized, testable assertion)
- Example (Bad): "The study investigates how exposure to blue light before sleep affects melatonin production." (Descriptive of method, not a general claim)
- **Be clear and concise**: Use straightforward language without unnecessary words.
- Example (Good): "The Eiffel Tower is in Paris."
- Example (Bad): "The Eiffel Tower, which is one of the most iconic landmarks in Europe and attracts millions of tourists every year, is located in Paris, France."
- **Exclude narrative context**: Focus on the factual assertion itself, not the surrounding story or background information.
- Example (Good): "Water boils at 100C under normal atmospheric pressure."
- Example (Bad): "In many cultures, people have believed for centuries that water boils at 100C, as scientists confirmed in the 18th century." (Includes unnecessary background information.)
- **Disregard all self-referential content**: Ignore any statements referring to the passage itself or the author intentions.
- Example (Good): "The Earth orbits the Sun."
- Example (Bad): "The study explains how the Earth orbits the Sun." (This is self-referential and refers to the passage itself.)
- **Be precise and objective**: Avoid ambiguity, subjective interpretation, or vague statements. Present claims as clear, verifiable facts.
- Example (Good): "The Great Wall of China stretches approximately 13,000 miles."
- Example (Bad): "The Great Wall of China is pretty long." (Vague and subjective.)
- **Be relevant to the broader debate or public discourse**: Focus on verification-worthy claims that introduce new information rather than merely restating common knowledge.
- Example (Good): "The global temperature has increased by about 1C since the late 19th century."
- Example (Bad): "The Earth is a planet." (Common knowledge and not contributing new, verifiable information.)
Present the output using the following format:
claim
claim
claim
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
PASSAGE: {text}
A.3 SciClaims claim extraction second step (refinement)
Now, using the information given by the passage, reformulate each individual claim to be fully understandable by itself, even without having the context from the passage or the rest of the claims from the list. Change the terminology or add context information to each claim if necessary.
A.4 SciClaims+CDP+CR (refinement)
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
Given a claim and the passage it was extracted from, reformulate the claim to fully adhere to **ALL** the following criteria.
- **Convey an insight, interpretation, or conclusion drawn from the passage that is testable and generalizable**: The claim should assert an outcome, capability, or effect rather than merely describing a method, aim, or process.
- Example (Good): "Neural networks outperform decision trees in image classification tasks." (Testable outcome)
- Example (Bad): "The proposed method aims to use neural networks for better image classification." (Descriptive, not assertive)
- **Be expressed as a meaningful, self-contained statement**: Each claim should be fully understandable on its own, without needing context from the passage or other claims. It must convey a complete, independent idea. If referencing a study, survey, result, or process, phrase it as a general, verifiable claim.
- Example (Good): "The Amazon rainforest is home to over 10 million species."
- Example (Bad): "As mentioned earlier, the Amazon is one of the most biodiverse places in the world." (Requires prior context and doesnt stand alone.)
- **Emphasize generalization and scientific assertion**: Avoid descriptive or narrative conclusions.
- Example (Good): "Exposure to blue light before sleep can reduce melatonin production." (Generalized, testable assertion)
- Example (Bad): "The study investigates how exposure to blue light before sleep affects melatonin production." (Descriptive of method, not a general claim)
- **Be clear and concise**: Use straightforward language without unnecessary words.
- Example (Good): "The Eiffel Tower is in Paris."
- Example (Bad): "The Eiffel Tower, which is one of the most iconic landmarks in Europe and attracts millions of tourists every year, is located in Paris, France."
- **Exclude narrative context**: Focus on the factual assertion itself, not the surrounding story or background information.
- Example (Good): "Water boils at 100C under normal atmospheric pressure."
- Example (Bad): "In many cultures, people have believed for centuries that water boils at 100C, as scientists confirmed in the 18th century." (Includes unnecessary background information.)
- **Disregard all self-referential content**: Ignore any statements referring to the passage itself or the author intentions.
- Example (Good): "The Earth orbits the Sun."
- Example (Bad): "The study explains how the Earth orbits the Sun." (This is self-referential and refers to the passage itself.)
- **Be precise and objective**: Avoid ambiguity, subjective interpretation, or vague statements. Present claims as clear, verifiable facts.
- Example (Good): "The Great Wall of China stretches approximately 13,000 miles."
- Example (Bad): "The Great Wall of China is pretty long." (Vague and subjective.)
- **Be relevant to the broader debate or public discourse**: Focus on verification-worthy claims that introduce new information rather than merely restating common knowledge.
- Example (Good): "The global temperature has increased by about 1C since the late 19th century."
- Example (Bad): "The Earth is a planet." (Common knowledge and not contributing new, verifiable information.)
Present the output using the following format:
{"original_claim": <str>, "refined_claim": <str>, "rationale": <str>}
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
CLAIM: {claim}
PASSAGE: {text}
A.5 SciClaims claim verification
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
You are a claim analyst. Upon receiving a claim and an evidence, your task is to figure out if the claim is either supported, contradicted or unrelated based exclusively on the evidence.
- If you are confident that the claim is supported by the evidence your answer will be "SUPPORT".
- If you are certain that the evidence directly contradicts the claim, your answer will be "CONTRADICT". Please note that if the claim is just not mentioned in the evidence, or if it is unrelated to the evidence, it does not mean it is contradicted. For that cases, the answer will be "NEI".
- If the evidence does not contain enough information or if it is not related to the claim, your answer will be "NEI", which stands for Not Enough Information.
Arrange the output as a JSON dictionary with the keys "response" and "evidence" and ensure your output is JSON-valid.
- The "response" values can only be "SUPPORT", "CONTRADICT" or "NEI".
- The "evidence" value must be a list of sentences from the evidence which are more related to your decision. If the decision is "NEI", this field will be empty.
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
CLAIM: {claim}
EVIDENCE: {evidence}
A.6 Phase 1 evaluation with judge model (Q1)
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
Given a sentence and a paragraph, answer the following question. Use exclusively the content of the paragraph to answer the question.
Is the sentence supported by the paragraph?
Return your response as a json dictionary, following this structure: {"answer":<Yes/No>, "rationale": <str>}
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
SENTENCE: {claim}
PARAGRAPH: {text}
A.7 Phase 1 evaluation with judge model (Q2-8)
Given a claim, answer the following question.
{QUESTION}
Return your response as a json dictionary, following this structure: {"answer":<Yes/No>, "rationale": <str>}
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
CLAIM: {claim}
A.8 Phase 2 evaluation with judge model
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
Given a claim and a paragraph, answer the following question.
Is the information contained in the paragraph useful to verify the claim?
Return your response as a json dictionary, following this structure: {"answer":<Yes/No>, "rationale": <str>}
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
CLAIM: {claim}
PARAGRAPH: {text}
A.9 Phase 3 evaluation with judge model
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
Given a claim and a paragraph, answer the following question.
Is the claim {SUPPORTED/REFUTED} by the paragraph?
Return your response as a json dictionary, following this structure: {"answer":<Yes/No>, "rationale": <str>}
<|begin_of_text|><|start_header_id|>user<|end_header_id|>
CLAIM: {claim}
PARAGRAPH: {text}
Appendix B Screenshots
Figure 2: Screenshot of the SciClaims Demo examples panel.
Figure 3: Screenshot of the SciClaims Demo results.
Appendix C Additional metrics
Method Rouge1 Rouge2 RougeL BertScore Similarity Score
Sentence tokenizer 0.3313 0.198 0.3030 0.8001 0.7211
Pan et al 0.278 0.1334 0.2555 0.8152 0.6561
Wright et al 0.2525 0.0762 0.222 0.862 0.6475
Noun Phrases Gen 0.2567 0.0953 0.2293 0.8753 0.6693
SciClaims 0.3387 0.1896 0.3084 0.8204 0.725
Table 7: Results in Claim Generation in SciFact
Method Rouge1 Rouge2 RougeL BertScore Similarity Score
Sentence tokenizer 0.2846 0.1332 0.268 0.8173 0.6549
Pan et al 0.3722 0.1801 0.3477 0.8816 0.7147
Wright et al 0.361 0.1559 0.3359 0.8761 0.7023
Noun Phrases Gen 0.3383 0.141 0.3146 0.8844 0.7018
SciClaims 0.3448 0.1722 0.3219 0.8244 0.7028
Table 8: Results in Claim Generation in FEVER
Method Rouge1 Rouge2 RougeL BertScore Similarity Score
Sentence tokenizer 0.2905 0.1333 0.2744 0.8587 0.6667
Pan et al 0.2796 0.134 0.2569 0.8155 0.7128
Wright et al 0.3375 0.1384 0.3188 0.8803 0.6786
Noun Phrases Gen 0.3213 0.1326 0.302 0.8881 0.6889
SciClaims 0.3533 0.1708 0.3345 0.8802 0.7184
Table 9: Results in Claim Generation in RoSE